Rico is twice as old as ryan. In four years rico's age will be 3 times ryan's age six years ago. What are there present age?
Answer by macston(5194)

x=Rico; y=Ryan
x=2y
x+4yrs=3(y-6yrs)Substitute for x.
2y+4yrs=3y-18yrs Add 18 to each side
2y+22yrs=3y Subtract 2y from each side.
22yrs=y ANSWER 1: Ryan is 22 years old
x=2y=2(22yrs)=44 yrs ANSWER 2: Rico is 44 years old.
CHECK:
x+4yrs=3(y-6yrs)
44+4yrs=3(22-6yrs)
48yrs=3(16yrs)
48yrs=48yrs